I was wondering how would you solve a linear system using subsitution? My teacher taught the class this in like 10 min and I'm not sure.. Can you show me to steps to answer -x+2y=-6 and 8x+y=31 thanks!
Found 2 solutions by stanbon, lynnlo:
Answer by stanbon(75887)

-x+2y=-6
8x+ y=31
-------------
You have a choice: solve the 1st equation for "x"
or solve the 2nd equation for "y":
-----
Solving for x you get:
x = 2y+6
----
Substitute for "x" in the 2nd equation and solve for "y":
8(2y+6)+y = 31
16y + 48 + y = 31
17y = -17
y = -1
-----------
Solve for "x" using x = 2y+6
x = 2*-1+6
x = 4
====================
Cheers,
Stan H.
====================

Answer by lynnlo(4176)

-X+2Y=-6
8X+Y=3
STEP (1)MULTIPLY FIRST EQUATION BY 8
AFTER YOU MULTIPLY YOU WILL HAVE
-8X+16Y=-48
8X+Y=3
STEP (2)ADD THE TWO EQUATIONS TOGETHER TO ELIMINATE X
17Y=-45
STEP(3) FIND Y
Y=-45/17
STEP (4)SUBSTITUTE THE VALUE FOR Y INTO THE ORIGINAL EQUATION TO SOLVE X
-X+2(-45/17)=-6
X=12/17
Y=-45/17